126

Pour en savoir plus

[AC93] Roberto M. Amadio and Luca Cardelli. Subtyping recursive types. Transactions on Programming Languages and Systems. ACM, 15(4) :575–631, 1993.

[AC94] Martín Abadi and Luca Cardelli. A theory of primitive objects : Untyped and first-order systems. In Theoretical Aspects of Computer Software, pages 296–320. Springer-Verlag, April 1994.

[AC95] Martín Abadi and Luca Cardelli. A theory of primitive objects : Second-order systems. Science of Computer Programming, 25(2-3) :81–116, December 1995. Preliminary version appeared in D. Sanella, editor, Proceedings of European Symposium on Programming, pages 1-24. Springer-Verlag, April 1994.

[BCP96] Kim Bruce, Luca Cardelli, and Benjamin Pierce. Comparing objects encodings. Presented at the FOOL’3 workshop. Available electronically in informal workshop proceedings, July 1996.

[CHC93] William R. Cook, Walter L. Hill, and Peter S. Canning. Inheritance is not subtyping. In Carl A. Gunter and John C. Mitchell, editors, Theoretical Aspects Of Object-Oriented Programming. Types, Semantics and Language Design. MIT Press, 1993.

[KP93] Dexter Kozen and Jens Palsberg. Efficient recursive subtyping. In Proc. 20th symp. Principles of Programming Languages, pages 419–428. ACM press, 1993.

[Oho90] Atsushi Ohori. Extending ML polymorphism to record structure. Technical Report CSC 90/R24, University of Glasgow, Department of Computer Science, September 1990.

[Rém94] Didier Rémy. Programming objects with ML-ART : An extension to ml with abstract and record types. In Masami Hagiya and John C. Mitchell, editors, Theoretical Aspects of Computer Software, volume 789 of Lecture Notes in Computer Science, pages 321–346. Springer-Verlag, April 1994.

[RV97] Didier Rémy and Jérôme Vouillon. Objective ML : A simple object-oriented extension to ML. In Proceedings of the 24th ACM Conference on Principles 128 POUR EN SAVOIR PLUS

of Programming Languages, January 1997. To appear. Also presented at the FOOL’3 workshop, available electronically in informal workshop proceedings.